# Beyond One World: Benchmarking Super Heros in Role-Playing Across Multiversal Contexts

### v1

@@ Abstract
Perapard Ngokpol1, Kun Kerdthaisong1, Pasin Buakhaw2,†, Pitikorn Khlaisamniang3, Supasate Vorathammathorn3, Piyalitt Ittichaiwong4,5,*, Nutchanon Yongsatianchot1,* 1Thammasat School of Engineering, Thammasat University 2Department of Computer Engineering and Digital Technology, Faculty of Engineering, Chulalongkorn University 3Artificial Intelligence Association of Thailand 4School of Biomedical Engineering & Imaging Sciences, King’s College London 5Siriraj Informatics and Data Innovation Center (SIData+), Faculty of Medicine, Siriraj Hospital, Mahidol University These authors contributed equally to this work. *Corresponding author
@@ 3.2 Dilemma Situations
To gauge whether an LLM can deliberate in line with a hero’s ethical compass, we created a bank of 1,080 multiple-choice dilemmas through a systematic generation and curation process. For each of the 90 character versions, we synthesized three distinct scenarios spanning four archetypal moral conflicts using Claude Sonnet 3.7 to correct grammar, then manually filtered them for canonical fidelity and linguistic clarity.

### v2

@@ Abstract
Perapard Ngokpol1thanks: Equal contribution., Kun Kerdthaisong [^1] footnotemark: 1, Pasin Buakhaw211footnotemark: 1,Pitikorn Khlaisamniang3, Supasate Vorathammathorn3,Piyalitt Ittichaiwong4,5thanks: Corresponding authors., Nutchanon Yongsatianchot122footnotemark: 2 1Faculty of Engineering, Thammasat School of Engineering, Thammasat University 2Department of Computer Engineering and Digital Technology, Faculty of Engineering, Chulalongkorn University 3Artificial Intelligence Association of Thailand 4School of Biomedical Engineering & Imaging Sciences, King’s College London 5Siriraj Informatics and Data Innovation Center (SIData+), Faculty of Medicine, Siriraj Hospital, Mahidol University
@@ 3.2 Dilemma Situations
To gauge whether an LLM can deliberate in line with a hero’s ethical compass, we created a bank of 1,080 multiple-choice dilemmas through a systematic generation and curation process. For each of the 90 character versions, we synthesized three distinct scenarios spanning four archetypal moral conflicts using Claude Sonnet 3.7, then manually filtered them for canonical fidelity and linguistic clarity.
@@ 5.4 Discussion
Acknowledgments
This research was supported by the Faculty of Engineering, Thammasat School of Engineering, Thammasat University also thanks to PreceptorAI that provides API for generate additional training data.